A sample of hydrogen atom is exposed to electromagnetic radiation of 0 1026A which causes
emission of induced radiation. Which of the following can be induced radiation?
(a) 0 6564A (b) 0 1216A (c) 0 1026A (d) 0 4863A

Grade:11

1 Answers

Vikas TU
7 years ago
Calculate energy first we get,
E = 12400/1026 = > 12.085 eV.
For H atom n =1 to n=2
delta E = 10.2
for next state it is 12.1
for next it is 12.75 .
Therfore the it would have the largest probabikity to absorb  with energy of 12.1 eV radiation.
Thus,
k = 12400/12.1 = > 1026.44 Angstorm